# A mixed finite element method for a class of fourth-order stochastic evolution equations with multiplicative noise

### v1

@@ Abstract.
We develop a fully discrete, semi-implicit mixed finite element method for approximating solutions to a class of fourth-order stochastic partial differential equations (SPDEs) with non-globally Lipschitz and non-monotone nonlinearities, perturbed by spatially smooth multiplicative Gaussian noise. The proposed scheme is applicable to a range of physically relevant nonlinear models, including the stochastic Landau–Lifshitz–Baryakhtar (sLLBar) equation, the stochastic convective Cahn–Hilliard equation with mass source, and the stochastic regularised Landau–Lifshitz–Bloch (sLLB) equation, among others. To overcome the difficulties posed by the interplay between the nonlinearities and the stochastic forcing, we adopt a ‘truncate-then-discretise’ strategy: the nonlinear term is first truncated before discretising the resulting modified problem. We show that the strong solution to the truncated problem converges in probability to that of the original problem. A fully discrete numerical scheme is then proposed for the truncated system, and we establish both convergence in probability and strong convergence (with quantitative rates) for the two fields used in the mixed formulation.
@@ 1. Introduction
$\boldsymbol{u}(0,\boldsymbol{x})=\boldsymbol{u}_{0}(\boldsymbol{x})\;$ for $\quad\text{for }\boldsymbol{x}\in\mathscr{D},$ (1.1c)
$\frac{\partial\boldsymbol{u}}{\partial\boldsymbol{n}}=\boldsymbol{0},\;\displaystyle{\frac{\partial\boldsymbol{H}}{\partial\boldsymbol{n}}=\boldsymbol{0}}\;$ for ($\quad\text{for }(t,\boldsymbol{x})\in(0,T)\times\partial\mathscr{D},$ (1.1d)
@@ 2.1. Notations
$\text{D}(\Delta):=\left\{\boldsymbol{v}\in\mathbb{H}^{2}:\frac{\partial\boldsymbol{v}}{\partial\boldsymbol{n}}=0\text{ on }\partial\mathscr{D}\right\}.$
If $X$ is a Banach space, $L^{p}(0,T;X)$ and $W^{k,p}(0,T;X)$ denote respectively the usual Lebesgue and Sobolev spaces of functions on $(0,T)$ taking values in $X$. The space $C([0,T];X)$ denotes the space of continuous functions on $[0,T]$ taking values in $X$. The space $L^{p}(\Omega;X)$ denotes the space of $X$-valued random variable with finite $p$-th moment, where $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space.
@@ 2.3. Existence, uniqueness, and regularity of solution
$\boldsymbol{u}\in L^{p}\big{(}\Omega;C^{\alpha}(0,T;\text{D}(A^{\beta}))\big{)}\cap L^{p}\big{(}\Omega;L^{2}(0,T;\text{D}(A))\big{)}.$
@@
$\boldsymbol{u}_{R}(0,\boldsymbol{x})=\boldsymbol{u}_{0}(\boldsymbol{x})\;$ for $\quad\text{for }\boldsymbol{x}\in\mathscr{D},$ (2.5c)
$\frac{\partial\boldsymbol{u}_{R}}{\partial\boldsymbol{n}}=\boldsymbol{0},\;\displaystyle{\frac{\partial\boldsymbol{H}_{R}}{\partial\boldsymbol{n}}=\boldsymbol{0}}\;$ for ($\quad\text{for }(t,\boldsymbol{x})\in(0,T)\times\partial\mathscr{D},$ (2.5d)
@@
$\boldsymbol{u}_{R}\in L^{p}\big{(}\Omega;C^{\alpha}(0,T;\text{D}(A^{\beta}))\big{)}\cap L^{p}\big{(}\Omega;L^{2}(0,T;\text{D}(A))\big{)}.$
@@
=[$=\mathbb{P}\left[\left\{\tau_{R}<T\right\}\cap\left\{\sigma_{R}=T\right\}\right]+\mathbb{P}\left[\left\{\tau_{R}<T\right\}\cap\left\{\sigma_{R}<T\right\}\right]$
@@ 2.4. Finite element approximation
Let $\mathcal{T}_{h}$ be a quasi-uniform triangulation of $\mathscr{D}\subset\mathbb{R}^{d}$ with maximal mesh-size $h$, and let $\mathbb{V}_{h}\subset\mathbb{W}^{1,\infty}$ be the Lagrange finite element space
@@
$\left\|{\boldsymbol{v}-\Pi_{h}\boldsymbol{v}}\right\|_{\mathbb{L}^{p}}+h\left\|{\nabla(\boldsymbol{v}-\Pi_{h}\boldsymbol{v})}\right\|_{\mathbb{L}^{p}}\leq Ch^{2}\left\|{\boldsymbol{v}}\right\|_{\mathbb{W}^{2,p}}.$ (2.11)
@@
$\left\|{\boldsymbol{v}-\mathcal{R}_{h}\boldsymbol{v}}\right\|_{\mathbb{L}^{p}}+h\left\|{\nabla(\boldsymbol{v}-\mathcal{R}_{h}\boldsymbol{v})}\right\|_{\mathbb{L}^{p}}$ $\leq Ch^{2}\left\|{\boldsymbol{v}}\right\|_{\mathbb{W}^{2,p}}.$ (2.13)
@@
Consequently, for any $p,q\in[1,\infty]$ such that $11/p+1/q=$ 1, by Hölder’s inequality we have
@@ 2.5. Identities and inequalities
    Suppose that $21/p+1/q=1/$ 2. For any $\boldsymbol{v}:\mathscr{D}\to\mathbb{R}^{3}$ and $\boldsymbol{w}:\mathscr{D}\to\mathbb{R}^{3}$,
@@
Now, let $p,q,r\in[1,\infty]$ be such that $11/p+1/q+1/r=$ 1. For each $\epsilon>0$, there exists a positive constant $C$ such that for any $\boldsymbol{v}\in\mathbb{W}^{1,q}\cap L^{p}$ and $\boldsymbol{w}\in\mathbb{L}^{r}$,
@@ 3. A fully-discrete mixed finite element method
[…]
@@
$\quad+\mathbb{E}\left[\sum_{j=1}^{n}k\left(\left\|{\nabla\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}+\left\|{\Delta_{h}\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)\left\|{\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2^{p}-2}\right]$
$\quad+\mathbb{E}\left[\left(k\sum_{j=1}^{n}\left\|{\nabla\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2^{p-1}}\right]+\mathbb{E}\left[\left(k\sum_{j=1}^{n}\left\|{\Delta_{h}\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2^{p-1}}\right]\leq C.$ (3.2)
@@
This completes the proof of inequality (3.2). ∎
@@
Let $(\boldsymbol{u},\boldsymbol{H})$ be the solution of (2.7) with initial data $\boldsymbol{u}_{0}\in\mathbb{H}^{2}$, and let $(\boldsymbol{u}_{h}^{n},\boldsymbol{H}_{h}^{n})$ be the solution to (3.1). Let $\Omega_{\kappa,m}$ be as defined in (3.32). Let $\boldsymbol{\theta}^{n}$ and $\boldsymbol{\xi}^{n}$ be as defined in (3.29) and (3.30), respectively. Assume that $h=O(k)$. Then for $n\in\{1,2,\ldots,N\}$, we have
$\mathbb{E}\left[\max_{m\leq n}\left(\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\boldsymbol{\theta}^{m}}\right\|_{\mathbb{L}^{2}}^{2}\right)\right]+k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left(\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}+\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right)\right]$ $\leq e^{C_{R,T}\kappa^{2}}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
$\quad-\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\big{(}\boldsymbol{u}(s)-\boldsymbol{u}_{h}^{\ell}\big{)}\times\boldsymbol{H}(s),\boldsymbol{\chi}_{h}\right\rangle\mathrm{d}s-\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}_{h}^{\ell}\times\big{(}\boldsymbol{H}(s)-\boldsymbol{H}_{h}^{\ell}\big{)},\boldsymbol{\chi}_{h}\right\rangle\mathrm{d}s$
@@
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle(\boldsymbol{\rho}^{\ell}+\boldsymbol{\theta}^{\ell}+\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell}))\times\boldsymbol{H}(s),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}_{h}^{\ell}\times(\boldsymbol{\eta}^{\ell}+\boldsymbol{\xi}^{\ell}+\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
@@
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle(\boldsymbol{\rho}^{\ell}+\boldsymbol{\theta}^{\ell}+\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell}))\times\boldsymbol{H}(s),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}_{h}^{\ell}\times(\boldsymbol{\eta}^{\ell}+\boldsymbol{\xi}^{\ell}+\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
@@
For the second term, by Young’s inequality, (2.13), and Proposition 2.2,
@@
Similarly, for the terms $I_{3}$ and $I_{4}$ we have
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck^{\frac{1}{2}-\delta},$ (3.40)
and noting (3.32),
$\left|{I_{4}}\right|$ $\leq Ch^{4}+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{H}(s)}\right\|_{\mathbb{L}^{2}}^{2}\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{4}}^{2}\mathrm{d}s\right]$
@@
$\leq Ch^{4}+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\sup_{s\in[t_{\ell-1},t_{\ell}]}\left\|{\boldsymbol{H}(s)}\right\|_{\mathbb{L}^{2}}^{2}\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{4}}^{2}\right]$
@@
$\leq Ch^{4}+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+C\kappa k^{1-\delta}.$
For the term $I_{5}$, by Hölder’s and Young’s inequalities and (2.22), for any $\epsilon>0$ we have
$\left|{I_{5}}\right|$ $\leq\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{4}}\left(\left\|{\boldsymbol{\eta}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}+\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}\right)\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\mathrm{d}s\right]$
$\leq Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{u}_{h}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}^{2}\right]+Ch^{4}+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
$\quad+\epsilon\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\|{\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})}\right\|_{\mathbb{L}^{2}}^{2}\mathrm{d}s\right]$
@@
$\leq Ch^{4}+Ck^{1-\delta}+C\kappa^{2}k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$

### v2

@@ Abstract.
We develop a fully discrete, semi-implicit mixed finite element method for approximating solutions to a class of fourth-order stochastic partial differential equations (SPDEs) with non-globally Lipschitz and non-monotone nonlinearities, perturbed by spatially smooth multiplicative Gaussian noise. The proposed scheme is applicable to a range of physically relevant nonlinear models, including the stochastic Landau–Lifshitz–Baryakhtar (sLLBar) equation, the stochastic convective Cahn–Hilliard equation with mass source, and the stochastic regularised Landau–Lifshitz–Bloch (sLLB) equation, among others. To overcome the difficulties posed by the interplay between the nonlinearities and the stochastic forcing, we adopt a ‘truncate-then-discretise’ strategy: the nonlinear term is first truncated before discretising the resulting modified problem. We show that the strong solution to the truncated problem converges in probability to that of the original problem. A fully discrete numerical scheme is then proposed for the truncated system, and we establish both convergence in probability and strong convergence (with quantitative rates) for the two fields used in the mixed formulation. Numerical simulations are provided to support the theoretical results.
@@ 1. Introduction
$\boldsymbol{u}(0,\boldsymbol{x})=\boldsymbol{u}_{0}(\boldsymbol{x})\;$ $\quad\text{for }\boldsymbol{x}\in\mathscr{D},$ (1.1c)
$\frac{\partial\boldsymbol{u}}{\partial\boldsymbol{n}}=\boldsymbol{0},\;\displaystyle{\frac{\partial\boldsymbol{H}}{\partial\boldsymbol{n}}=\boldsymbol{0}}\;$ $\quad\text{for }(t,\boldsymbol{x})\in(0,T)\times\partial\mathscr{D},$ (1.1d)
@@ 2.1. Notations
If $X$ is a Banach space, $L^{p}(0,T;X)$ and $W^{k,p}(0,T;X)$ denote respectively the usual Lebesgue and Sobolev spaces of strongly measurable functions on $(0,T)$ taking values in $X$. The space $L^{p}(\Omega;X)$ denotes the space of strongly measurable $X$-valued random variable with finite $p$-th moment, where $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The space $\mathcal{C}^{0}([0,T];X)$ and $\mathcal{C}^{\alpha}([0,T];X)$ denote respectively the space of continuous functions and the space of $\alpha$-Hölder continuous functions on $[0,T]$ taking values in $X$. For brevity, we will denote the spaces $L^{p}(0,T;X)$, $\mathcal{C}^{0}([0,T];X)$, and $\mathcal{C}^{\alpha}([0,T];X)$ by $L^{p}_{T}(X)$, $\mathcal{C}^{0}_{T}(X)$, and $\mathcal{C}^{\alpha}_{T}(X)$, respectively.
Throughout this paper, we denote the scalar product in a Hilbert space $H$ by $\langle\cdot,\cdot\rangle_{H}$ and its corresponding norm by $\|\cdot\|_{H}$. The expectation of a random variable $Y$ will be denoted by $\mathbb{E}[Y]$. We will not distinguish between the scalar product of $\mathbb{L}^{2}$ vector-valued functions taking values in $\mathbb{R}^{3}$ and the scalar product of $\mathbb{L}^{2}$ matrix-valued functions taking values in $\mathbb{R}^{3\times 3}$, and denote them by $\langle\cdot,\cdot\rangle$.
@@ 2.3. Existence, uniqueness, and regularity of solution
$\boldsymbol{u}\in L^{p}\big(\Omega;\mathcal{C}^{\alpha}(0,T;\text{D}(A^{\beta}))\big)\cap L^{p}\big(\Omega;L^{2}(0,T;\text{D}(A))\big).$
@@
$\boldsymbol{u}_{R}(0,\boldsymbol{x})=\boldsymbol{u}_{0}(\boldsymbol{x})\;$ $\quad\text{for }\boldsymbol{x}\in\mathscr{D},$ (2.5c)
$\frac{\partial\boldsymbol{u}_{R}}{\partial\boldsymbol{n}}=\boldsymbol{0},\;\displaystyle{\frac{\partial\boldsymbol{H}_{R}}{\partial\boldsymbol{n}}=\boldsymbol{0}}\;$ $\quad\text{for }(t,\boldsymbol{x})\in(0,T)\times\partial\mathscr{D},$ (2.5d)
@@
$\boldsymbol{u}_{R}\in L^{p}\big(\Omega;\mathcal{C}^{\alpha}(0,T;\text{D}(A^{\beta}))\big)\cap L^{p}\big(\Omega;L^{2}(0,T;\text{D}(A))\big).$
@@
$=\mathbb{P}\left[\left\{\tau_{R}<T\right\}\cap\left\{\sigma_{R}=T\right\}\right]+\mathbb{P}\left[\left\{\tau_{R}<T\right\}\cap\left\{\sigma_{R}<T\right\}\right]$
@@ 2.4. Finite element approximation
Let $\{\mathcal{T}_{h}\}_{h>0}$ be a family of quasi-uniform triangulations of $\mathscr{D}\subset\mathbb{R}^{d}$ with maximal mesh-size $h$, and let $\mathbb{V}_{h}\subset\mathbb{W}^{1,\infty}$ be the Lagrange finite element space
@@
$\left\|{\boldsymbol{v}-\Pi_{h}\boldsymbol{v}}\right\|_{\mathbb{L}^{p}}+h\left\|{\nabla(\boldsymbol{v}-\Pi_{h}\boldsymbol{v})}\right\|_{\mathbb{L}^{p}}\leq Ch^{s}\left\|{\boldsymbol{v}}\right\|_{\mathbb{W}^{s,p}},\quad s\in\{1,2\}.$ (2.11)
@@
$\left\|{\boldsymbol{v}-\mathcal{R}_{h}\boldsymbol{v}}\right\|_{\mathbb{L}^{p}}+h\left\|{\nabla(\boldsymbol{v}-\mathcal{R}_{h}\boldsymbol{v})}\right\|_{\mathbb{L}^{p}}$ $\leq Ch^{s}\left\|{\boldsymbol{v}}\right\|_{\mathbb{W}^{s,p}},\quad s\in\{1,2\}.$ (2.13)
@@
Consequently, for any $p,q\in[1,\infty]$ such that $1/p+1/q=1$, by Hölder’s inequality we have
@@ 2.5. Identities and inequalities
    Suppose that $1/p+1/q=1/2$. For any $\boldsymbol{v}:\mathscr{D}\to\mathbb{R}^{3}$ and $\boldsymbol{w}:\mathscr{D}\to\mathbb{R}^{3}$,
@@
Now, let $p,q,r\in[1,\infty]$ be such that $1/p+1/q+1/r=1$. For each $\epsilon>0$, there exists a positive constant $C$ such that for any $\boldsymbol{v}\in\mathbb{W}^{1,q}\cap L^{p}$ and $\boldsymbol{w}\in\mathbb{L}^{r}$,
@@ 3. A fully-discrete mixed finite element method
$\left\{\begin{aligned} \left\langle\boldsymbol{u}_{h}^{n}-\boldsymbol{u}_{h}^{n-1},\boldsymbol{\chi}_{h}\right\rangle&=k\left\langle\boldsymbol{H}_{h}^{n},\boldsymbol{\chi}_{h}\right\rangle+k\left\langle\nabla\boldsymbol{H}_{h}^{n},\nabla\boldsymbol{\chi}_{h}\right\rangle-k\gamma\left\langle\boldsymbol{u}_{h}^{n}\times\boldsymbol{H}_{h}^{n},\boldsymbol{\chi}_{h}\right\rangle+k\left\langle\mathcal{C}(\boldsymbol{u}_{h}^{n}),\boldsymbol{\chi}_{h}\right\rangle\\ &\quad+k\left\langle\mathcal{M}_{R}(\boldsymbol{u}_{h}^{n-1}),\boldsymbol{\chi}_{h}\right\rangle+\left\langle G(\boldsymbol{u}_{h}^{n-1}),\boldsymbol{\chi}_{h}\right\rangle\overline{\Delta}W^{n},\\ \left\langle\boldsymbol{H}_{h}^{n},\boldsymbol{\phi}_{h}\right\rangle&=-\left\langle\nabla\boldsymbol{u}_{h}^{n},\nabla\boldsymbol{\phi}_{h}\right\rangle+\left\langle f_{R}(\boldsymbol{u}_{h}^{n-1}),\boldsymbol{\phi}_{h}\right\rangle,\end{aligned}\right$. (3.1)
@@
$\quad+\mathbb{E}\left[\sum_{j=1}^{n}k\left(\left\|{\nabla\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}+\left\|{\Delta_{h}\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)\left\|{\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2^{p}-2}\right]+\mathbb{E}\left[\left(k\sum_{j=1}^{n}\left\|{\nabla\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2^{p-1}}\right]$
$\quad+\mathbb{E}\left[\left(k\sum_{j=1}^{n}\left\|{\Delta_{h}\boldsymbol{u}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2^{p-1}}\right]+{\mathbb{E}\left[\left(k\sum_{j=1}^{n}\left\|{\boldsymbol{H}_{h}^{j}}\right\|_{\mathbb{L}^{2}}^{2}\right)^{2^{p-1}}\right]}\leq C.$ (3.2)
@@
This completes the proof of inequality (3.2), upon noting the definition of $\boldsymbol{H}_{h}^{n}$ in the second equation of (3.1). ∎
@@
Let $(\boldsymbol{u},\boldsymbol{H})$ be the solution of (2.7) with initial data $\boldsymbol{u}_{0}\in\mathrm{D}(A^{\frac{1}{2}})$, and let $(\boldsymbol{u}_{h}^{n},\boldsymbol{H}_{h}^{n})$ be the solution to (3.1). Let $\Omega_{\kappa,m}$ be as defined in (3.32). Let $\boldsymbol{\theta}^{n}$ and $\boldsymbol{\xi}^{n}$ be as defined in (3.29) and (3.30), respectively. Assume that $h=O(k)$. Then for $n\in\{1,2,\ldots,N\}$, we have
$\mathbb{E}\left[\max_{m\leq n}\left(\mathds{1}_{\Omega_{\kappa,m-1}}\left\|{\boldsymbol{\theta}^{m}}\right\|_{\mathbb{L}^{2}}^{2}\right)\right]+k\sum_{\ell=1}^{n}\mathbb{E}\left[\mathds{1}_{\Omega_{\kappa,\ell-1}}\left(\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}+\left\|{\boldsymbol{\xi}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right)\right]$ $\leq e^{C_{R,T}\kappa}\left(h^{2}+k^{\frac{1}{2}-\delta}\right),$
@@
$\quad-\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\big(\boldsymbol{u}(s)-\boldsymbol{u}_{h}^{\ell}\big)\times\boldsymbol{H}_{h}^{\ell},\boldsymbol{\chi}_{h}\right\rangle\mathrm{d}s-\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}(s)\times\big(\boldsymbol{H}(s)-\boldsymbol{H}_{h}^{\ell}\big),\boldsymbol{\chi}_{h}\right\rangle\mathrm{d}s$
@@
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle(\boldsymbol{\rho}^{\ell}+\boldsymbol{\theta}^{\ell}+\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell}))\times\boldsymbol{H}_{h}^{\ell},\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
$\quad-\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}(s)\times(\boldsymbol{\eta}^{\ell}+\boldsymbol{\xi}^{\ell}+\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]$
@@
$\quad-{\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle(\boldsymbol{\rho}^{\ell}+\boldsymbol{\theta}^{\ell}+\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell}))\times\boldsymbol{H}_{h}^{\ell},\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]}$
$\quad-{\mathbb{E}\left[\max_{m\leq n}\sum_{\ell=1}^{m}\mathds{1}_{\Omega_{\kappa,\ell-1}}\int_{t_{\ell-1}}^{t_{\ell}}\left\langle\boldsymbol{u}(s)\times(\boldsymbol{\eta}^{\ell}+\boldsymbol{\xi}^{\ell}+\boldsymbol{H}(s)-\boldsymbol{H}(t_{\ell})),\boldsymbol{\theta}^{\ell}\right\rangle\mathrm{d}s\right]}$
@@
For the second term, by Young’s inequality, (2.13), and the Hölder continuity in time of the solution given by Proposition 2.2,
@@
Similarly, for the term $I_{3}$, we have
@@
$\leq\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck^{\frac{1}{2}-\delta}.$ (3.40)
Next, we estimate $I_{4}$. By Hölder’s and Young’s inequalities, we obtain
@@
$\left|{I_{4}}\right|$ $\leq Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left(\left\|{\boldsymbol{\rho}^{\ell}}\right\|_{\mathbb{L}^{4}}+\left\|{\boldsymbol{u}(s)-\boldsymbol{u}(t_{\ell})}\right\|_{\mathbb{L}^{4}}\right)\left\|{\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{4}}\right]$
@@
$\leq Ckh^{2}\mathbb{E}\left[\sum_{\ell=1}^{n}\left\|{\boldsymbol{u}}\right\|_{L^{\infty}_{T}(\mathbb{W}^{1,4})}^{2}\left\|{\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$
@@
$\quad+Ck\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}k^{1-\delta}\left\|{\boldsymbol{u}}\right\|_{\mathcal{C}^{1/2-\delta}_{T}(\mathbb{H}^{2})}^{2}\left\|{\boldsymbol{H}_{h}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]+\epsilon k\mathbb{E}\left[\sum_{\ell=1}^{n}\mathds{1}_{\Omega_{\kappa,\ell-1}}\left\|{\nabla\boldsymbol{\theta}^{\ell}}\right\|_{\mathbb{L}^{2}}^{2}\right]$